RFBG Commissioning with Adaptive Secondary - 26 March 2010 UT

Night 7:

Observer: DMiller, JHeidt, FBian (LBTO)
Telescope Operator: DHuerta
Software:
Instrument Support:: TShih
Telescope Support: JUrban
Adaptive Support: GBrusa (LBTO)

Summary

Night started OK, wind increased as function of night, lost the last 3hrs due to high winds

Got a good set of field aberration data (40" tilt)

Did also some transformation data, but not of sufficient quality

Lots of problems with collimation, coma & astigmatism at beginning of night, not clear where it comes from

Preparations

Single frame names refer to: irtc.20100326.xxxxxx.fits

Cube names refer to: irtc.20100326_00xxx_cube.fit

Plan for the night: Do a large set of transformation data, collect field aberration measurements

If seeing is bad: take smaller set of transformation data using Dougs IDL-script to test it

Details

Open

01:45 Open, some thicker clouds in north + some Cirrus , T_out = 0.5C, T_Mirror = 1.7C, T_Shell = ~7.2C (reference body), Hum = 30%, windy

Initial Pointing and Collimation Adjustment

Each night in this RFBG commissioning run, we go to a bright star (Pollux) at Elevation 80 degrees. Then we adjust the vertex tip-tilt (RX/RY) of the secondary to get rid of the coma that we see in the extrafocal pupil on the IRTC.

01:55 pointto Pollux gs=0, TelM = track, PAMode=no, no star, adjust to yesterday values (CA=-42, IE=-41), EL = 78, OK, star there, adjust *CA=-17 ,IE=-48*
DX M2 Global X = 2.5, Y = -2.63, Z = 1.9, RX = -500, RY = 0 , not too bad

02:12 pointto M5_1210 gs=0, TelM = Track , EL = 26, AZ = 170, at very-low elevation to adjust AZ-zeropoint, pointing OK, star close by, Adjust IA = -260, keep IE=-48, CA = -24. Some coma, Doug adjusts global offsets to:
DX M2 Global X = 2.5, Y = -1.43, Z = 1.9, RX = -600, RY = 0.

This follows the trend we saw on the left side of the telescope. IA is near -390 arcsec if the temperature has warmed by 5 degC, and IA is near -270 if the temperature has cooled.

02:26 pointto M5_0354, GS=0, TelM=track, check for pointing near Stone-G, OK, perfect!

02:29 pointto BS9143, GS=0, TelM=active, check for collimation/seeing, wind at secondary 10m/s, need to close all vent doors, now its about 2m/s, RMS@WFE below 400 but WFS image looks ugly

Dougs has put collimation-images in the dir /OldRepository/AGW_Data/100326/: images 1 - 11

02:56 pointto M5_0354 again, GS=0, TelM=track, to get rid of strong Coma, no idea where it came from, David cleared active optics
DX M2 Global X = 2.5, Y = -1.5, Z = 2.3, RX = -580, RY = 0, CA=-30, IE=-52.

03:09 pointo BS9143, GS=0, TelM=guide, still bad, lots of Coma

03:16 pointto M5_0354 again, GS=0, TelM=track, clear active optics again, extrafocal image shows strong coma

03:25 pointto BS9126, GS=0, TelM=guide, EL=79, also Coma albeit high elevation

03:30 pointto BS9126, GS=0, TelM=acquire, EL=79,
DX M2 Global X = 2.5, Y = -4.3, Z = 1.7, RX = -300, RY = 0, CA=-40, IE=-64.

03:39 pointto BS9126, GS=0, TelM=guide, Doug tries now to collimate using his IDL-routines, image 17 in dir given above, seems to work, ask Guido to reapply a flat on secondary to see if it improves WFS-camera images, done, looks MUCH better (see image 57), image quality as god as 0.7"

04:16 pointto BS9143, GS=0, TelM=active, check for collimation via GCS (and not Dougs-routines), lots of astigmatism, Guido flattens secondary again, no improvement ???

04:27 pointto BS9148, GS=0, TelM=active, check for collimation again (astigmatism) similar EL as for BS9126 where it worked, astigmatism present, but went away with active optics correction

04:43 pointto BS9143, GS=0, TelM=active, check for collimation again (Zernikes are not reported to PSF), astigmatism, went away once Zernikes are reported to PSF

04:57 pointto StoneGagw, GS=0, TelM=active, let collimate, OK, moon is unfort. close (9h, +15deg, vs. 10h, 0deg for StoneG), but lets try one set of transform data (also as check for Dougs routines)

05:07 cl < tran_StoneG_mix1.cl, guider reports mostly below 0.8"

PA = 0 transtar (26) 050910 cube 158
transtar (16) 051207 cube 159
transtar (24) 051341 cube 160

PA = 30 transtar (25) 051747 cube 161
transtar (19) 052028 cube 162
transtar (23) 052309 cube 163, guiding failed, not useful
transtar (13) 052548 cube 164, guiding failed, not useful

PA = 90 transtar (20) 052831 cube 165
transtar (13) 053114 cube 166
transtar (10) 053356 cube 167

PA = 135 transtar (13) 053650 cube 168, guiding failed, not useful
transtar (10) 053924 cube 169
transtar (7) 054209 cube 170

PA = 180 transtar (10) 054453 cube 171
transtar (7) 054734 cube 172, guiding probably not good
transtar (4) 055016 cube 173

PA = 210 transtar (11) 055321 cube 174, no guidestar, not useful
transtar (8) 055554 cube 175
transtar (5) 055839 cube 176

PA = 270 transtar (4) 060141 cube 177, no guidestar, not useful
transtar (3) 060437 cube 178, no guidestar, not useful
moon is too close, doesn't makes sense to continue

06:20 Problem with some actuators of secondary, need to wait a little bit

06:57 pointto BS9145, GS=0, TelM=active, check collimation, secondary seems to be flat now, no idea what the shape is, astigmatism, but collimates OK, guider reports ~0.9"

07:04 pointto StoneIagw, GS=0, TelM=active, let collimate, OK, got E-mail by Andrew asking for field aberration data

Field Aberration Measurements on BS9152
Data are in /OldRepository/AGW_Data/100326_Field

07:09 pointto BS9152, GS=0, TelM=active, let collimate, OK

07:29 pointto BS9152, GS=0, TelM=acquire, PAMode=no, centering star by David,

07:36 Change DX M2 global to X = 2.5, Y = -3.63, Z = 2.08, RX = -520, RY = 0, clear active optics from M2, than collimate, change Rx, Ry to adjust out coma, did a "tilt" of "+40" on OPEM2, total collimation for M2 X=0.457, Y=-3.63, Z=2.08, RX=-520, RY=-3.92.799, RZ=0, restart GCS (twice)

DX M1 total X=-0.741, Y=0.346, Z=-0.361, RX=-18.323, RY=21.609.

08:04 pointto BS9152, GS=0, TelM=active, let collimate, OK, guider reports ~0.6"

08:13 pointto BS9152 gs=0, TelM=guide, OK centered

GStar wfscimage#
Sequence 1
0 1 - 2
9 3 - 5
0 6 - 7
17 8 - 10
0 11 - 16
18 18 - 20
0 21 - 23
7 24 - 26
0 27 - 28
4 29 - 31
0 32 - 33
1 34 - 36
0 37 - 38
Sequence 2
9 39 - 41
0 42 - 43
17 44 - 46
0 47 - 48
18 49 - 51
0 52 - 53
7 54 - 56
0 57 - 58
4 59 - 61
0 62 - 64
1 65 - 67
0 68 - 72
Sequence 3
9 73 - 75
0 76 - 77
17 78 - 80
0 81 - 82
18 83 - 85
0 86 - 87
7 88 - 90

Notes: Preset after image 56: GCS disconnected from AGW while slewing, restart OSS
Images 68-72, more than required but sent erroneously a preset too early
Wind has increased considerably during the last 2 hrs,

10:17 Dome closed due to wind exceeding 20 m/s, sequence 3 could not be finished

11:02 Removed the 40" of tilt that we had applied to the secondary

Wind was too strong the rest of night, T_out=-2.2, T_Mirror=-1.0, Hum=20%, Wind > 20 m/s, sky is clear

-- JochenHeidt - 26 Mar 2010
Topic revision: r13 - 30 Mar 2010, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback